OpenStudy (anonymous):

PLEASE HELP Find the slope of the straight line and it's y intercept y + 5 = 10

OpenStudy (anonymous):

do you mean: y+5x=10

OpenStudy (anonymous):

y=mx+b where m is the slope and b is the y intercept.

OpenStudy (mertsj):

\[y+5=10\] \[y=5\] horizontal line, slope = 0

OpenStudy (anonymous):

yeah, you r going ringht.